### v1

@@ 3.1 Multi-star Fitting in EXOFASTv2
As discussed in §2.4, a stellar companion to TOI-4462 was detected 0.4 away, which was blended in all catalog photometry and a significant factor in the dilution of the transit light curves. Given that the probability of a chance alignment is low, and the high Gaia Re-normalized Unit Weight Error (RUWE) of 3.13, we assumed this companion is bound to the primary star. We undid the deblending that SPOC applies to the TESS lightcurves so that we could more accurately model it based on our multi-component SED model. We modeled both stars simultaneously, each with their own MIST evolutionary model, while assuming that the age, initial metallicity, distance, and extinction is the same for both stars. In addition, we modeled a spectral energy distribution for each star, constraining the sum of both stars with the catalog photometry of the unresolved TOI-4462 system, and the difference between the two stars with the AO photometry from PHARO, as shown in Figure 1 (Left). The dilution of the transit light curve was accounted for by fitting for dilution terms that were then constrained by the multi-component SED model, integrated at the transit-observed bands assuming a 5% floor in the theoretical dilution from the model atmospheres. That is, we applied an adaptive prior penalty of
@@ 4 Discussion
The 5 BD-mass companions presented here increase the population of transiting brown dwarfs to over 50. While this number is expected to continue growing, it is worth analyzing the sizable population that has been put together thus far in the context of planet-like and star-like formation. We may find that early trends or features observed in the growing population from previous efforts have been reinforced, or lost significance (perhaps even disappearing) in the wake of new discoveries. One such feature of particular interest is the so-called ”brown dwarf desert” and its potential role in dividing the brown dwarfs into distinct planet-like and star-like groups. We discuss these trends below in §4.1 and §4.2. We also note that one of the BD-hosting stars presented here, TOI-5882, has a significant absorption feature at 6708 Å which we attribute to lithium. We discuss in §4.3 the implications of this, as well as how it affects our determination of the system’s age in.
@@ 4.2 Transiting Brown Dwarf Metallicities
Eccentricity is likely not the only parameter that could offer insight into at which critical companion mass the dominant formation mechanism turns over from planet-like to star-like. For nearly two decades we have known of the giant planet-metallicity correlation, in which hot Jupiter hosting stars tend to be more metal rich than their counterparts with no discovered planets (Gonzalez 1997; Santos et al. 2003; Fischer & Valenti 2005). If the low-mass transiting BDs are predominantly forming in the same way as the hot Jupiters, then we should expect their host stars to exhibit the same metallicity enhancement when compared to the high-mass BD hosts. Schlaufman (2018) tested this hypothesis using the metallicities of transiting companions in the range to show that transition between core accretion and fragmentation may be as low as . However, at the time of this study, there were only 27 transiting companions known between limiting the ability to probe potential higher-mass transition points. Now that we have access to significantly more systems in this mass regime, we can better probe the same 42 transition. In Figure 4 we show a preliminary look at testing this hypothesis. Qualitatively, it appears that the lower-mass companions ( ) preferentially orbit more metal-rich host stars. However, a two-sample Kolmogovorov-Smirnov test yields a -value of 0.35, too high to reject the null hypothesis that the high mass and low mass samples are drawn from the same underlying distribution.
We note also a few important caveats for interpreting the metallicity distribution. First, we chose a 42 cutoff for historical reasons in order to compare to the original hypothesis presented by Ma & Ge (2014) as well as the eccentricity distribution presented in §4.1. It may be that a more appropriate boundary separating the population will be found after a more comprehensive analysis which may be in better agreement with the lower-mass Schlaufman (2018) transition. With just 26 companions below 42 in our sample, we chose note to investigate possible lower-mass transitions. A more in-depth analysis including the population of giant planets will need to be done to more precisely probe a lower mass transition. An unbiased sample of the population of companions across the substellar limit will also be required to better combat the selection effects currently affecting this population. We note also that the metallicities presented here are the reported values from each system’s discovery which have been measured using a variety of different techniques and therefore the underlying biases affecting each measurement are not explored here.
@@ 4.3 Lithium Detected in TOI-5882
The presence of lithium in stars is typically interpreted as an indicator of youth. This is due to the temperatures and pressures in the core being sufficiently high to destroy Li, which results in Li visible on the stellar surface slowly depleting as transport occurs between the core and the surface of the star (Soderblom et al. 2014). Despite this, we claim that TOI-5882 is likely not a young star, since we found no other signs of youth. Specifically, we found no evidence of rapid rotation, comoving companions, or an infrared excess. To verify this, we performed a period search on each sector of TOI-5882’s TESS light curves to characterize the rotation of the host star. Since young stars are typically born rapidly rotating and gradually spin down over time, an age can often be inferred from a star’s rotation period if it is below the Kraft break (Bouma et al. 2023). We found a significant peak in the periodogram at 9.6 days, however, we are hesitant to adopt this as the true rotation period since periodicity beyond 1/3 of a TESS observing sector ( 9 days) can be unreliable due aliases induced by the TESS observing strategy and processing of light curves. Nevertheless, this rotation period would still be much too long to be consistent with a young age given this star’s effective temperature (Bouma et al. 2023).
@@
We also searched for nearby comoving stars using FriendFinder 666https://github.com/adamkraus/comove (Tofflemire et al. 2021) since their presence would indicate that TOI-5882 and its hypothetical nearby comovers have not yet dispersed from their birth location and hence would be young. FriendFinder identifies all nearby sources that fall within a selected search radius, and calculates the predicted tangential velocity for each source, assuming that they have Galactic velocity components (,,) identical to TOI-5882. This predicted is then compared to the true which is derived from Gaia proper motions. Using a physical search radius of 30 pc around TOI-5882 and a difference between the predicted and measured of km s-1 we find no evidence that TOI-5882 is part of a comoving group.
@@
Finally, we looked for an infrared excess since young stellar objects that retain a circumstellar disk show increased emission at infrared wavelengths (Cotten & Song 2016). We rule out an infrared excess for TOI-5882 via our SED fitting in the global analysis where we see no significant infrared emission above the blackbody model in any of the WISE , , and bandpasses. While the lack of these additional youth indicators do not completely rule out the possibility of a young host star, we believe it is more likely that TOI-5882 is a late main-sequence star, as indicated by our most probable EXOFASTv2 solution. This older age could then imply that the presence of Li is due to the infall of planetary material onto the host star. For a deeper dive into the origin of Li in TOI-5882, including its potential as system that has undergone a planetary engulfment, we alert the reader to Kotten et al. (in prep).
@@ 5 Conclusions
In this paper, we present the discovery of 11 new transiting companions from the TESS mission. We collected photometric time-series, spectroscopic, and high resolution imaging follow-up as a part of the TESS Follow-up Observing Program (TFOP) to rule our false positives and further characterize each system. Using EXOFASTv2, we performed a global fit on each system using the space and ground-based transits, spectroscopic RVs, and archival photometry to characterize both the host stars, and their transiting companions. We found that 5 of these systems are brown dwarfs () and 6 of them are very low mass stars from . This contribution to the transiting brown dwarf population increases it to 54 systems, a milestone that represents the population outgrowing burdens of small sample statistics.
@@
Finally, we noted the presence of Li in the spectrum of TOI-5882, the host star of our lowest mass BD. We measured the equivalent width of the Li line and search for other signs of youth. Seeing no evidence of rapid rotation, nearby comoving stars, or infrared excess, we adopt the age provided by our global EXOFASTv2 fit and do not interpret the presence of Li as a sign of youth. Instead we noted that the Li may actually be a signature of engulfed planetary material, and that more work will be required to explore this hypothesis.
@@ Acknowledgements
This paper includes data collected by the TESS mission that are publicly available from the Mikulski Archive for Space Telescopes (MAST). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This research has made use of the NASA Exoplanet Archive and the Exoplanet Follow-up Observation Program (ExoFOP; DOI: 10.26134/ExoFOP5) website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. We acknowledge financial support from the Agencia Estatal de Investigación of the Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033 and the ERDF “A way of making Europe” through project PID2021-125627OB-C32, and from the Centre of Excellence “Severo Ochoa” award to the Instituto de Astrofisica de Canarias. This work makes use of observations from the LCOGT network. Part of the LCOGT telescope time was granted by NOIRLab through the Mid-Scale Innovations Program (MSIP). MSIP is funded by NSF. This paper is based on observations made with the Las Cumbres Observatory’s education network telescopes that were upgraded through generous support from the Gordon and Betty Moore Foundation. Based on observations obtained at the Hale Telescope, Palomar Observatory, as part of a collaborative agreement between the Caltech Optical Observatories and the Jet Propulsion Laboratory operated by Caltech for NASA. DRC and CAC acknowledge partial support from NASA Grant 18-2XRP18_2-0007. Based in part on observations obtained at the Southern Astrophysical Research (SOAR) telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the U.S. National Science Foundation NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). This work is partly supported by JSPS KAKENHI Grant Number JP24H00017, JP24K00689 and JSPS Bilateral Program Number JPJSBP120249910. This paper is based on observations made with the MuSCAT2 instrument, developed by ABC, at Telescopio Carlos Sánchez operated on the island of Tenerife by the IAC in the Spanish Observatorio del Teide.
@@
NV is supported by the NASA FINESST program. The postdoctoral fellowship of KB is funded by F.R.S.-FNRS grant T.0109.20 and by the Francqui Foundation. KAC and CNW acknowledge support from the TESS mission via subaward s3449 from MIT. I.A.S. acknowledges the support of M.V. Lomonosov Moscow State University Program of Development. F. M. acknowledge support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) through grant PID2023-152906NA-I00. LM acknowledges financial contribution from PRIN MUR 2022 project 2022J4H55R.

### v2

@@ 3.1 Multi-star Fitting in EXOFASTv2
As discussed in §2.4, a stellar companion to TOI-4462 was detected 0.4 away, which was blended in all catalog photometry and a significant factor in the dilution of the transit light curves. Given that the probability of a chance alignment is low, and the high Gaia Re-normalized Unit Weight Error (RUWE) of 3.13, we assumed this companion is bound to the primary star. We undid the deblending that SPOC applies to the TESS lightcurves so that we could more accurately model it based on our multi-component SED model. We modeled both stars simultaneously, each with their own MIST evolutionary model, while assuming that the age, initial metallicity, distance, and extinction is the same for both stars. In addition, we modeled a spectral energy distribution for each star, constraining the sum of both stars with the catalog photometry of the unresolved TOI-4462 system, and the difference between the two stars with the AO photometry from PHARO, as shown in Figure 1 (Left). We therefore fit for dilution terms that were then constrained by the multi-component SED model, integrated at the transit-observed bands assuming a 5% floor in the theoretical dilution from the model atmospheres. That is, we applied an adaptive prior penalty of
@@ 4 Discussion
The 5 BD-mass companions presented here increase the population of transiting brown dwarfs to over 50. While this number is expected to continue growing, it is worth analyzing the sizable population that has been put together thus far in the context of planet-like and star-like formation. We have also added 6 new transiting low-mass stars to the population . Accurate mass and radius measurements are rare for these objects which will be vital for anchoring our understanding of stellar formation and evolution. We may find that early trends or features observed in the growing population from previous efforts have been reinforced, or lost significance (perhaps even disappearing) in the wake of new discoveries. One such feature of particular interest is the so-called ”brown dwarf desert” and its potential role in dividing the brown dwarfs into distinct planet-like and star-like groups. We discuss these trends below in §4.1 and §4.2. We also note that one of the BD-hosting stars presented here, TOI-5882, has a significant absorption feature at 6708 Å which we attribute to lithium. We discuss in §4.3 the implications of this, as well as how it affects our determination of the system’s age.
@@ 4.2 Transiting Brown Dwarf Metallicities
Eccentricity is likely not the only parameter that could offer insight into at which critical companion mass the dominant formation mechanism turns over from planet-like to star-like. For nearly three decades we have known of the giant planet-metallicity correlation, in which hot Jupiter hosting stars tend to be more metal rich than their counterparts with no discovered planets (Gonzalez 1997; Santos et al. 2003; Fischer & Valenti 2005). If the low-mass transiting BDs are predominantly forming in the same way as the hot Jupiters, then we should expect their host stars to exhibit the same metallicity enhancement when compared to the high-mass BD hosts. Schlaufman (2018) tested this hypothesis using the metallicities of transiting companions in the range to show that transition between core accretion and fragmentation may be as low as . However, at the time of this study, there were only 27 transiting companions known between limiting the ability to probe potential higher-mass transition points. Now that we have access to significantly more systems in this mass regime, we can better probe the same 42 transition. In Figure 4 we show a preliminary look at testing this hypothesis. Qualitatively, it appears that the lower-mass companions ( ) preferentially orbit more metal-rich host stars. However, a two-sample Kolmogovorov-Smirnov test yields a -value of 0.35, too high to reject the null hypothesis that the high mass and low mass samples are drawn from the same underlying distribution.
We note also a few important caveats for interpreting the metallicity distribution. First, we chose a 42 cutoff for historical reasons in order to compare to the original hypothesis presented by Ma & Ge (2014) as well as the eccentricity distribution presented in §4.1. It may be that a more appropriate boundary separating the population will be found after a more comprehensive analysis which may be in better agreement with the lower-mass Schlaufman (2018) transition. With just 26 companions below 42 in our sample, we chose note to investigate possible lower-mass transitions. A more in-depth analysis including the population of giant planets will need to be done to more precisely probe a lower mass transition. An unbiased sample of the population of companions across the substellar limit will also be required to better understand the selection effects currently affecting this population. We note also that the metallicities presented here are the reported values from each system’s discovery which have been measured using a variety of different techniques and therefore the underlying biases affecting each measurement are not explored here.
@@ 4.3 Lithium Detected in TOI-5882
The presence of lithium in stars is typically interpreted as an indicator of youth. This is due to the temperatures and pressures in the core being sufficiently high to destroy Li, which results in Li visible on the stellar surface slowly depleting as transport occurs between the core and the surface of the star (Soderblom et al. 2014). Despite this, we claim that TOI-5882 is likely not a young star, since we found no other signs of youth. To verify this, we performed a period search on each sector of TOI-5882’s TESS light curves to characterize the rotation of the host star. Since young stars are typically born rapidly rotating and gradually spin down over time, an age can often be inferred from a star’s rotation period if it is below the Kraft break (Bouma et al. 2023). We found a significant peak in the periodogram at 9.6 days, however, we are hesitant to adopt this as the true rotation period since periodicity beyond 1/3 of a TESS observing sector ( 9 days) can be unreliable due to aliases induced by the TESS observing strategy and processing of light curves. Even if we were to believe that the 9.6 day periodicity is truly due to stellar rotation, it is still anomalous when compared to the observed rotation periods of young stars. T Tauri stars for example rarely exhibit rotation periods longer than 8 days (Serna et al. 2021), and gyrochronology shows that a 9.6 day period would be indicative of an age of approximately 1 Gyr given this star’s effective temperature (Bouma et al. 2023). While a young age for TOI-5882 cannot be conclusively ruled out by its rotation, it is unlikely, especially combined with the lack of other youth indicators.
@@
For example, we also searched for nearby comoving stars using FriendFinder 666https://github.com/adamkraus/comove (Tofflemire et al. 2021) since their presence would indicate that TOI-5882 and its hypothetical nearby comovers have not yet dispersed from their birth location and hence would be young. FriendFinder identifies all nearby sources that fall within a selected search radius, and calculates the predicted tangential velocity for each source, assuming that they have Galactic velocity components (,,) identical to TOI-5882. This predicted is then compared to the true which is derived from Gaia proper motions. Using a physical search radius of 30 pc around TOI-5882 and a difference between the predicted and measured of km s-1 we find no evidence that TOI-5882 is part of a comoving group. Furthermore, the nearest star-forming regions in Cygnus, where TOI-5882 is located, are much further away ( kpc; Reipurth & Schneider 2008).
@@
Finally, we looked for an infrared excess as well as H emission. Young stellar objects that retain a circumstellar disk show increased emission at infrared wavelengths (Cotten & Song 2016). We ruled out an infrared excess for TOI-5882 via our SED fitting in the global analysis where we see no significant infrared emission above the blackbody model in any of the WISE , , and bandpasses. The presence of H lines in emission is also characteristic of active young stars (Briceño et al. 2019), and we found no evidence of such emission. While the lack of these additional youth indicators do not completely rule out the possibility of a young host star, we believe it is more likely that TOI-5882 is a late subgiant star, as indicated by our most probable EXOFASTv2 solution. This older age could then imply that the presence of Li is due to the infall of planetary material onto the host star. For a deeper dive into the origin of Li in TOI-5882, including its potential as system that has undergone a planetary engulfment, we alert the reader to Kotten et al. (in prep).
@@ 5 Conclusions
In this paper, we present the discovery of 11 new transiting companions from the TESS mission. We collected photometric time-series, spectroscopic, and high resolution imaging follow-up as a part of the TESS Follow-up Observing Program (TFOP) to rule our false positives and further characterize each system. Using EXOFASTv2, we performed a global fit on each system using the space and ground-based transits, spectroscopic RVs, and archival photometry to characterize both the host stars, and their transiting companions. We found that 5 of these systems are brown dwarfs () and 6 of them are very low mass stars from . This contribution to the transiting brown dwarf population increases it to 54 systems, a milestone that represents the population outgrowing the burdens of small sample statistics.
@@
Finally, we noted the presence of Li in the spectrum of TOI-5882, the host star of our lowest mass BD. We measured the equivalent width of the Li line and search for other signs of youth. Seeing no evidence of youth from any of the other indicators that we examined, we adopt the age provided by our global EXOFASTv2 fit and do not interpret the presence of Li as a sign of youth. Instead, we noted that the Li may actually be a signature of engulfed planetary material, and that more work will be required to explore this hypothesis.
@@ Acknowledgements
This paper includes data collected by the TESS mission that are publicly available from the Mikulski Archive for Space Telescopes (MAST; STScI 2018). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. Resources supporting this work were provided by the NASA High-End Computing (HEC) Program through the NASA Advanced Supercomputing (NAS) Division at Ames Research Center for the production of the SPOC data products. This research has made use of the NASA Exoplanet Archive and the Exoplanet Follow-up Observation Program (ExoFOP; DOI: 10.26134/ExoFOP5) website, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. We acknowledge financial support from the Agencia Estatal de Investigación of the Ministerio de Ciencia e Innovación MCIN/AEI/10.13039/501100011033 and the ERDF “A way of making Europe” through project PID2021-125627OB-C32, and from the Centre of Excellence “Severo Ochoa” award to the Instituto de Astrofisica de Canarias. This work makes use of observations from the LCOGT network. Part of the LCOGT telescope time was granted by NOIRLab through the Mid-Scale Innovations Program (MSIP). MSIP is funded by NSF. This paper is based on observations made with the Las Cumbres Observatory’s education network telescopes that were upgraded through generous support from the Gordon and Betty Moore Foundation. Based on observations obtained at the Hale Telescope, Palomar Observatory, as part of a collaborative agreement between the Caltech Optical Observatories and the Jet Propulsion Laboratory operated by Caltech for NASA. DRC and CAC acknowledge partial support from NASA Grant 18-2XRP18_2-0007. Based in part on observations obtained at the Southern Astrophysical Research (SOAR) telescope, which is a joint project of the Ministério da Ciência, Tecnologia e Inovações (MCTI/LNA) do Brasil, the U.S. National Science Foundation NOIRLab, the University of North Carolina at Chapel Hill (UNC), and Michigan State University (MSU). This work is partly supported by JSPS KAKENHI Grant Number JP24H00017, JP24K00689 and JSPS Bilateral Program Number JPJSBP120249910. This paper is based on observations made with the MuSCAT2 instrument, developed by ABC, at Telescopio Carlos Sánchez operated on the island of Tenerife by the IAC in the Spanish Observatorio del Teide. We thank the staff of the Observatoire de Haute-Provence for their support at the 1.93 m telescope and on SOPHIE.
@@
NV is supported by the NASA FINESST program. The postdoctoral fellowship of KB is funded by F.R.S.-FNRS grant T.0109.20 and by the Francqui Foundation. KAC and CNW acknowledge support from the TESS mission via subaward s3449 from MIT. I.A.S. acknowledges the support of M.V. Lomonosov Moscow State University Program of Development. F. M. acknowledge support from the Agencia Estatal de Investigación del Ministerio de Ciencia, Innovación y Universidades (MCIU/AEI) through grant PID2023-152906NA-I00. LM acknowledges financial contribution from PRIN MUR 2022 project 2022J4H55R. TF’s work is supported by the French National Research Agency in the framework of the Investisse-ments d’Avenir program (ANR-15-IDEX-02), through the funding of the “Origin of Life” project of the Grenoble-Alpes University